# MambaVO: Deep Visual Odometry Based on Sequential Matching Refinement and Training Smoothing

### v1

@@ Abstract
Deep visual odometry has demonstrated great advancements by learning-to-optimize technology. This approach heavily relies on the visual matching across frames. However, ambiguous matching in challenging scenarios leads to significant errors in geometric modeling and bundle adjustment optimization, which undermines the accuracy and robustness of pose estimation. To address this challenge, this paper proposes MambaVO, which conducts robust initialization, Mamba-based sequential matching refinement, and smoothed training to enhance the matching quality and improve the pose estimation in deep visual odometry. Specifically, when a new frame is received, it is matched with the closest keyframe in the maintained Point-Frame Graph (PFG) via the semi-dense based Geometric Initialization Module (GIM). Then the initialized PFG is processed by a proposed Geometric Mamba Module (GMM), which exploits the matching features to refine the overall inter-frame pixel-to-pixel matching. The refined PFG is finally processed by deep BA to optimize the poses and the map. To deal with the gradient variance, a Trending-Aware Penalty (TAP) is proposed to smooth training by balancing the pose loss and the matching loss to enhance convergence and stability. A loop closure module is finally applied to enable MambaVO++. On public benchmarks, MambaVO and MambaVO++ demonstrate SOTA accuracy performance, while ensuring real-time running performance with low GPU memory requirement. Codes will be publicly available.
@@ 1 Introduction
Visual Odometry (VO) is the task of tracking an agent’s six-DOF poses when interacting with the physical world, which is a critical capability for robots, self-driving cars and other autonomous agents. Early works mainly rely on geometric feature methods to build VO systems[37, 3, 40], in which, the frontend extracts feature points and conducts feature matching[42, 30, 56], while the backend optimizes poses and build maps using filtering[16] or optimization methods[26, 9]. However, the geometric features are less effective in challenging environments, such as scenes with weak textures.
Building on the progress in computer vision and deep learning, early approaches have explored using neural networks to directly regress poses [52, 34, 29, 44, 54]. However, these methods suffer from poor accuracy and lack generalization to unseen environments. A more promising paradigm is the concept of learning to optimize [46, 48, 50], where neural networks are utilized to extract features and perform matching, while differentiable nonlinear optimization layers [46, 39] refine poses based on geometric residuals. This type of method has achieved state-of-the-art results in the field of deep visual odometry[50, 33].
@@
To tackle the issue, we propose MambaVO, a novel VO system that refines inter-frame image matching and produces more accurate and robust poses using a Mamba-based module. MambaVO introduces three key components. First, a Point-Frame Graph (PFG) captures observation relationships, enabling robust initialization and sequence-based matching refinement. Second, the Geometric Initialization Module (GIM) uses semi-dense matching [55] and PnP [27, 60] to predict pixel correspondences and initialize poses, extracting features for fusion in subsequent matching refinement. Third, the Geometric Mamba Module (GMM) refines pixel matches by leveraging historical tokens and fusion features through Mamba blocks, adjusting pixel correspondences and frame weights within the PFG. To ensure stable training, we further introduce the Trending-Aware Penalty (TAP), which balances pose and matching losses, addressing gradient variance issues [21]. We also extend MambaVO to MambaVO++, incorporating loop closure for global optimization in SLAM. MambaVO and MambaVO++ achieve state-of-the-art accuracy with real-time performance and reduced GPU memory usage across multiple benchmarks.
@@ 2.2 Learning to Optimize for Visual Odometry
2.3 State Space Model
@@ 3.1 Geometric Initialization Module: GIM
After receiving the $t$-th image, the nearest keyframe $r$ in PFG serves as the reference frame for initial matching and feature extraction. We utilize Dino-v2 [38] to extract context features, and EfficientLoFTR [55] provides $k$ initially matched coordinates $\{[p_{r}^{i},p_{t}^{i}]\}_{i=1}^{k}$ and pixel-wise geometric features $\mathbf{G}_{t}=\{\mathbf{G}_{t}^{i}\}_{i=1}^{k}$ from the fine-matching module in EfficientLoFTR. These $k$ matches are used to estimate the initial pose $\mathbf{T}_{t}$ of the frame $t$. We apply Metric3D [22] to estimate the depth for each matched point in the reference frame, denoted by $\{d_{r}^{i}\}_{i=1}^{k}$. By obtaining the 3D points and by the 3D-2D correspondences, PnP (Perspective-n-Point) [27] calculates the initial pose $\mathbf{T}_{t}$ and refine the 3D coordinates of the matched points:
@@ 3.1.2 Matching Feature Preparation
We prepare matching features for matching refinement by the Geometric Mamba Module (Sec. 3.2). The matched pixels in frame $i$ are used to query the context features from Divo-v2 encoder, yielding the pixel-wise context features $\mathbf{C}_{t}=\{\mathbf{C}_{t}^{i}\}_{i=1}^{k}$. Context features $\mathbf{C}_{t}$ and geometric features $\mathbf{G}_{t}$ are fused via concatenation, 1D convolution, and ReLU activation to obtain the fused matching feature $\mathbf{F}_{t}^{i}$ for each matched pixel in the pixel fusion process:
@@ 3.2.1 History Fusion
Odometry is essentially a time series problem, and historical information affects current inter-frame matching through sequential pose transformation and overlapping areas. We maintain the history token $H_{t-1}\in\mathbb{R}^{k\times 384}$ to represent the historical information and fuse with $\mathbf{F}_{t}$ in History Fusion. Specifically, we perform cross-attention between $\mathbf{F}_{t}^{i}$ and $H_{t-1}$ on each edge in PFG:
@@ 5 Experiments
We evaluate MambaVO and MambaVO++ on four datasets: EuRoC [1], TUM-RGBD [45], KITTI [15], and TartanAir [53]. Our model is trained only on the TartanAir training set, without retraining or fine-tuning on other datasets, which is the same as [50, 21, 48]. We train our network for 196k steps with a batch size of 1. Training takes 2.5 days on one RTX 3090 GPU.
@@ 5.1 Analysis on Pose Estimation
As the results in Tab. 1-Tab. 4, our method performs robustly across all scenarios, while previous methods fail on certain sequences (marked “✗”). MambaVO and MambaVO++ achieve the SOTA results (on AVG column) in both loop closure settings. Notably, MambaVO reduces the pose estimation error by 19%-22% on average comparing to DROID-VO [48], DPVO [50], and V2V [21] on indoor datasets like EuRoC, TUM-RGBD, and TartanAir.
@@ 5.4 Time and Memory
MambaVO can run in real-time on a single RTX 3090 GPU. All of the results are obtained under the settings shown in the Tab. 2, Tab. 3, Tab. 4.
We also compared GPU memory usage for inference with current state-of-the-art deep visual odometry methods. Experimental results demonstrate that our method has smaller GPU memory consumption, as the results shown in Fig. 6.
@@ 6 Conclusion
This paper proposes MambaVO, a novel deep visual odometry system to address the accuracy and robustness of deep VO based on learning to optimize. We conduct reliable initialization, sequential matching refinement, and training smoothing based on the Mamba architecture. Our method achieves the state-of-the-art results on TartanAir, EuRoC, TUM-RGBD, and KITTI.
Limitations: Our proposed MambaVO can only obtain sparse maps because only semi-dense matching pixels in GIM are used to build the map. In the future, we will utilize dense information and 3D Gaussian Splatting [25] technology to obtain high-fidelity maps for rich scene representation while maintaining localization accuracy.
Appendix
Appendix A Implementation Details
A.1 Geometric Initialization Module
We use pretrained EfficientLoFTR [55] in the Geometric Initialization Module (GIM) to obtain initial matches. EfficientLoFTR performs both coarse and fine-level matching. The fine-level stage provides pixel-level correspondences along with their features, which serve as our initial matches and geometric features. Since semi-dense matching can produce numerous matching pixels per frame, we randomly select 100 matches per frame for subsequent pose estimation.
We employ Dino-v2 [38] to generate generalized and robust semantic features. Specifically, we adopt the small version with registers as the pretrained weights of Dino-v2 model. For each input frame, Dino-v2 produces patch-level features with 384 channels. These patch features are reorganized and deconvolved to the pixel level to obtain semantic features for each pixel. Using the initial matching coordinates, the corresponding semantic features are queried and fused with geometric features to form the matching feature (Sec. 3.1). The resulting matching feature is then fed into the Geometric Mamba Module.
A.2 Geometric Mamba Module
Figure 7: Illustration of Mamba blocks in Geometric Mamba Module.
In the Geometric Mamba Module, we set the number of Mamba blocks $B$ to 3. The input to Mamba at frame $t$ consists of the history tokens from frame $t-1$ and all matching tokens of frames $t-\mathcal{W}$ to $t$, in chronological order. Mamba blocks output updated history tokens and matching tokens.
This design of input leverages Mamba’s long-distance modeling capabilities, utilizing historical interactions to adjust matching tokens and estimate the refinement and weight of each matching pixel.
Appendix B Qualitative Results
B.1 Matching Comparison
The accuracy of visual odometry heavily depends on computing visual correspondences between frames. In MambaVO, we first use semi-dense matching for initial correspondences, then refine them using the Mamba architecture with temporal information. Experimental results demonstrate that our matching approach outperforms previous visual odometry methods in terms of pose accuracy between frames (Sec. 5.2).
To intuitively demonstrate the matching quality of MambaVO, we visualize the matching pixels of MambaVO and the previous SOTA method DPVO [50] in Fig. 8. DPVO initializes 96 patches for matching per frame. During visualization, we remove invalid matches (those beyond the image coordinate range) and mark only the center points for clarity.
Compared to DPVO, which often produces ambiguous, inaccurate matches, MambaVO delivers more accurate, consistent, and evenly distributed matching points on the image plane, benefiting subsequent pose estimation.
Figure 8: Matching results of two consecutive frames in DPVO and MambaVO. The matching quality of MambaVO, including accuracy and distribution, is better than DPVO [50], which is the previous SOTA deep visual odometry.
B.2 Trajectory and Map Visualization
We visualize the results of MambaVO++ on EuRoC [1], TartanAir [53], KITTI [15] and TUM-RGBD [45] and the trajectories are illustrated in Fig. 9 and Fig. 10 respectively. The sparse maps are visualized in Fig. 11.
Among the four datasets, TartanAir is a simulation dataset, while the other three are collected from real-world environments. TartanAir covers both indoor and outdoor scenes, with trajectories ranging from a few meters to several tens of meters. The EuRoC dataset, collected by drones in indoor environments, shows that our method’s results align closely with the ground truth in most sequences. TUM-RGBD is a challenging indoor dataset characterized by erratic camera motion and significant motion blur. As seen in the experimental results (Tab. 4) and trajectory plots (Fig. 10), our method remains effective under these demanding conditions. KITTI, an outdoor autonomous driving dataset, typically features large-scale, high-speed movements. To evaluate our method’s performance on long-distance scenarios, we selected five sequences, including both those with and without loop closures, as in Fig. 10.
Figure 9: Qualitative visualization of MambaVO++ on EuRoC and TartanAir. The blue line represents the trajectory estimated by MambaVO++, and the red line represents the ground truth. The results show that our estimated trajectory almost completely coincides with the ground truth.
Figure 10: Qualitative visualization of MambaVO++ on KITTI and TUM-RGBD. The blue line represents the trajectory estimated by MambaVO++, and the red line represents the ground truth. The results show that our estimated trajectory almost completely coincides with the ground truth.
Figure 11: The proposed MambaVO can produce trajectories and sparse maps. We visualize the trajectories and sparse maps of KITTI (a and b) and EuRoC (c and d) respectively. The blue line represents the trajectories and the points respresent the map points.

### v2

@@ Abstract
Deep visual odometry has demonstrated great advancements by learning-to-optimize technology. This approach heavily relies on the visual matching across frames. However, ambiguous matching in challenging scenarios leads to significant errors in geometric modeling and bundle adjustment optimization, which undermines the accuracy and robustness of pose estimation. To address this challenge, this paper proposes MambaVO, which conducts robust initialization, Mamba-based sequential matching refinement, and smoothed training to enhance the matching quality and improve the pose estimation. Specifically, the new frame is matched with the closest keyframe in the maintained Point-Frame Graph (PFG) via the semi-dense based Geometric Initialization Module (GIM). Then the initialized PFG is processed by a proposed Geometric Mamba Module (GMM), which exploits the matching features to refine the overall inter-frame matching. The refined PFG is finally processed by differentiable BA to optimize the poses and the map. To deal with the gradient variance, a Trending-Aware Penalty (TAP) is proposed to smooth training and enhance convergence and stability. A loop closure module is finally applied to enable MambaVO++. On public benchmarks, MambaVO and MambaVO++ demonstrate SOTA performance, while ensuring real-time running.
@@ 1 Introduction
Visual Odometry (VO) is the task of tracking an agent’s six-DOF poses when interacting with the physical world, which is critical for robots, self-driving cars and other autonomous agents [23, 55, 54]. Early works mainly rely on geometric feature methods [38, 3, 41], in which, the frontend extracts feature points and conducts feature matching [43, 31, 60], while the backend optimizes poses and build maps using filtering [16] or optimization methods [27, 9, 56]. However, the geometric features are less effective in challenging environments, such as scenes with weak textures.
Building on the progress in computer vision and deep learning, early approaches have explored using neural networks to directly regress poses [53, 35, 30, 45]. However, these methods suffer from poor accuracy and lack generalization to unseen environments. A more promising paradigm is the concept of learning to optimize [47, 49, 51], where neural networks are utilized to extract features and perform matching, while differentiable nonlinear optimization layers [47, 40] refine poses based on geometric residuals. This type of method has achieved state-of-the-art results in the field of deep visual odometry[51, 34].
@@
To tackle the issue, we propose MambaVO, a novel VO system that refines inter-frame image matching and produces more accurate and robust poses using a Mamba-based module. MambaVO introduces three key components. First, a Point-Frame Graph (PFG) captures observation relationships, enabling robust initialization and sequence-based matching refinement. Second, the Geometric Initialization Module (GIM) uses semi-dense matching [59] and PnP [28, 64] to predict pixel correspondences and initialize poses, extracting features for fusion in subsequent matching refinement. Third, the Geometric Mamba Module (GMM) refines pixel matches by leveraging historical tokens and fusion features through Mamba blocks, adjusting pixel correspondences and frame weights within the PFG. To ensure stable training, we further introduce the Trending-Aware Penalty (TAP), which balances pose and matching losses, addressing gradient variance issues [21]. We also extend MambaVO to MambaVO++, incorporating loop closure for global optimization in SLAM.
@@ 2.2 Learning to Optimize for Visual Odometry
2.3 State Space Models
@@ 3.1 Geometric Initialization Module: GIM
After receiving the $t$-th image, the nearest keyframe $r$ in time order in PFG serves as the reference frame for initial matching and feature extraction. We utilize Dino-v2 [39] to extract context features, and EfficientLoFTR [59] provides $k$ initially matched coordinates $\{[p_{r}^{i},p_{t}^{i}]\}_{i=1}^{k}$ and pixel-wise geometric features $\mathbf{G}_{t}=\{\mathbf{G}_{t}^{i}\}_{i=1}^{k}$ from the fine-matching module in EfficientLoFTR. These $k$ matches are used to estimate the initial pose $\mathbf{T}_{t}$ of the frame $t$. We apply Metric3D [22] to estimate the depth for each matched point in the reference frame, denoted by $\{d_{r}^{i}\}_{i=1}^{k}$. By obtaining the 3D points and by the 3D-2D correspondences, PnP (Perspective-n-Point) [28] calculates the initial pose $\mathbf{T}_{t}$ and refine the 3D coordinates of the matched points:
@@ 3.1.2 Matching Feature Preparation
As in Fig. 2, we prepare matching features for matching refinement by the Geometric Mamba Module (Sec. 3.2). The matched pixels in frame $i$ are used to query the context features from Divo-v2 encoder, yielding the pixel-wise context features $\mathbf{C}_{t}=\{\mathbf{C}_{t}^{i}\}_{i=1}^{k}$. Context features $\mathbf{C}_{t}$ and geometric features $\mathbf{G}_{t}$ are fused via concatenation, 1D convolution, and ReLU activation to obtain the fused matching feature $\mathbf{F}_{t}^{i}$ for each matched pixel in the pixel fusion process:
@@ 3.2.1 History Fusion
Odometry is essentially a time series problem, and historical information affects current inter-frame matching through sequential pose transformation and overlapping areas. We maintain the history token $H_{t-1}\in\mathbb{R}^{k\times 384}$ to represent the historical information and fuse with $\mathbf{F}_{t}$ in History Fusion. The history token for the first frame is initialized from the mathing features in GIM. Specifically, we perform cross-attention between $\mathbf{F}_{t}^{i}$ and $H_{t-1}$ on each edge in PFG:
@@ 5 Experiments
We evaluate MambaVO and MambaVO++ on four datasets: EuRoC [1], TUM-RGBD [46], KITTI [15], and TartanAir [57]. Our model is trained only on the TartanAir training set, without retraining or fine-tuning on other datasets, which is the same as [51, 21, 49]. We train our network for 196k steps with a batch size of 1. Training takes 2.5 days on one RTX 4090 GPU.
@@ 5.1 Analysis on Pose Estimation
As the results in Tab. 1-Tab. 4, our method performs robustly across all scenarios, while previous methods fail on certain sequences (marked “✗”). Our Methods achieve the SOTA results (on AVG column) in both loop closure settings. Notably, MambaVO reduces the pose estimation error by 19%-22% on average comparing to DROID-VO [49], DPVO [51], and V2V [21] on indoor datasets like EuRoC, TUM-RGBD, and TartanAir.
@@ 5.4 Time and Memory
MambaVO can run in real-time on a single RTX 3090 GPU. On EuRoC, we achieve an average real-time performance of 22Hz, exceeding the camera frame rate of 20Hz. On the TUM-RGBD and KITTI datasets, the frame rates are 30Hz and 23Hz, respectively. All of the results are obtained under the settings shown in the Tab. 2, Tab. 3, Tab. 4.
We also compared GPU memory usage for inference with current state-of-the-art deep visual odometry methods. Experimental results demonstrate that our method has smaller GPU memory consumption during inference, as the results shown in Fig. 6.
@@ 6 Conclusion
This paper proposes MambaVO, a novel deep visual odometry system to address the accuracy and robustness of deep VO based on learning to optimize. We conduct reliable initialization, sequential matching refinement, and training smoothing based on the Mamba architecture. Our method achieves the state-of-the-art results on TartanAir, EuRoC, TUM-RGBD, and KITTI. In the future, we will utilize dense information and 3D Gaussian Splatting [26] technology to obtain high-fidelity maps for rich scene representation while aintaining localization accuracy.